OpenStudy (anonymous):

Which set of equations would be used to solve this word problem? An airplane flew 3 hours with a 30 mph head wind. The return trip with a tail wind of the same speed took 2 hours. Find the speed of the plane in still air. d = 3(r − 30) and d = 2(r + 30) d = 3(r + 30) and d = 2(r − 30) d = 5(r + 30)

OpenStudy (anonymous):

3 (r - 30) = 2 (r + 30) r = 150